A corporation issue a new bond at 5% interest rate compounded annually that pay 10000 at maturity five years later. Only one 10000 bullet payment in five years later. What is the purchase price for the bond at an issue date?
1
Expert's answer
2020-10-13T11:17:49-0400

P.V "= 10000 \times \frac{1}{(1+r)^{n}}"

P.V. "= 10000 \times \frac{1}{(1+0.05)^{5}}"

P.V. "= \frac{10000}{1.27628} = 7827"